OpenStudy (anonymous):

how does (sin(x))^2(cos(x))^2=(1/4)(sin(2x)^2

OpenStudy (anonymous):

use identities to make them equal

OpenStudy (anonymous):

inversin(2)/2

OpenStudy (anonymous):

sin(2x) = 2sin(x)cos(x) 4sin^(2)(x)cos^(2)(x)

OpenStudy (anonymous):

sorry thought i saw a + on the left

OpenStudy (anonymous):

(1/4)(sin(2x)^2 = sin(2x)sin(2x) = 4sin^(2)(x)cos^(2)(x)/4 = sin^(2)(x)cos^(2)(x)

OpenStudy (anonymous):

there easy

OpenStudy (anonymous):

got it!